OpenStudy (anonymous):

needing help on algebra 1....

OpenStudy (just_one_last_goodbye):

@Brainybeauty

OpenStudy (just_one_last_goodbye):

hope she can help you ^_^ god luck

OpenStudy (just_one_last_goodbye):

good luck*

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 1.Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

Kevin where's your question?

OpenStudy (just_one_last_goodbye):

@Brainybeauty i meant can u help her please

OpenStudy (anonymous):

Ohh!!! Okay sure! Why not?

OpenStudy (just_one_last_goodbye):

cool i owe u ^_^

OpenStudy (anonymous):

See, the formula for exponential growth is a(r^n -1) / r-1

OpenStudy (anonymous):

yess..

OpenStudy (anonymous):

Here, a is the principal population and r is the number by which it is multiplied each year. n is the number of years.

OpenStudy (anonymous):

So the formula will be \[f(x) = x(r ^{n}-1) \div r-1\] That will be the total number of snails at the end of each year.

OpenStudy (anonymous):

(r-1) is in brackets

OpenStudy (anonymous):

so what do i do with that formula? do i solve it?

OpenStudy (anonymous):

You're supposed to find the formula, right?

OpenStudy (anonymous):

im not sure. in # 1 it says to create my own exponetial function.

OpenStudy (anonymous):

for example, f(x) = 200 (1.03)2^x but that function is in number two so it cant be used. and im not sure how to come put with my own.

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

f(x) = 200 (1.03)^2x

OpenStudy (anonymous):

**

OpenStudy (anonymous):

Okay, for example the number of snails is 200. And every year the number of snails doubles each year. So the formula for 15 years will be --- \[200(2^{15} - 1) \div (2-1)\]

OpenStudy (anonymous):

okay. thanks. im still kinda confused but ill manage ..

OpenStudy (anonymous):

Okay I'll try to explain again oaky?

OpenStudy (anonymous):

The principal pop. is x The rate is r*100 per cent

OpenStudy (just_one_last_goodbye):

@Ayeeitshannaahhh since brainly helped can u give her a medal?

OpenStudy (anonymous):

okay

OpenStudy (just_one_last_goodbye):

thanks Hannah ^_^